OpenStudy (bloomlocke367):

The water level varies from 12 inches at low tide to 52 inches at high tide. Low tide occurs at 9:15 a.m. and high tide occurs at 3:30 p.m. What is a cosine function that models the variation in inches above and below the water level as a function of time in hours since 9:15 a.m.?

OpenStudy (bloomlocke367):

@phi @amistre64 @Preetha

OpenStudy (bloomlocke367):

@rational @SolomonZelman @Preetha @confluxepic

OpenStudy (bloomlocke367):

Can one of you help?

OpenStudy (amistre64):

start with cosine function and adjust it

OpenStudy (bloomlocke367):

how? @amistre64

OpenStudy (amistre64):

what does the information tell us?

OpenStudy (amistre64):

or rather, what are the parts of a cosine function? how do we write a generic cosine function?

OpenStudy (bloomlocke367):

that the min is 12 and the max is 52, right?

OpenStudy (amistre64):

good, and the midline is between them

OpenStudy (amistre64):

what is the distance from 12 to 52?

OpenStudy (amistre64):

strategy define the most general generic cosine function there is. use the information in the problem to build that function into the perfect model of the problem.

OpenStudy (bloomlocke367):

the distance is 40, so the amplitude is 20

OpenStudy (bloomlocke367):

@amistre64

OpenStudy (amistre64):

using words instead of letters, the generic cos function is Amp cos(per(t-phase)) + mid you found amp, what is midline?

OpenStudy (bloomlocke367):

and the most generic cosine function is y=cos(x), right?

OpenStudy (amistre64):

no, cos(x) is too specific, there are many more parts to a trig function then x

OpenStudy (bloomlocke367):

what do you mean?

OpenStudy (amistre64):

i mean what i said

OpenStudy (amistre64):

a completly generic cosine function is: Amp cos(per(t-phase)) + mid why????

OpenStudy (bloomlocke367):

what's (t-phase)?

OpenStudy (amistre64):

t is the independant input, and phase is the shifting of the function left or right

OpenStudy (amistre64):

do all cosine function have the value cos(0) = 1?

OpenStudy (bloomlocke367):

yes?

OpenStudy (amistre64):

really cos(t-pi/2), when t = 0 is equal to 1?

OpenStudy (amistre64):

might have misasked the first time.

OpenStudy (amistre64):

can functions be shifted left and right?

OpenStudy (bloomlocke367):

yes, they can be

OpenStudy (amistre64):

how do we shift a function f(x) ?? what gets adjusted?

OpenStudy (bloomlocke367):

it depends on the function, right? like a quadratic function shifts if the vertex changes

OpenStudy (amistre64):

it doesnt depend on the function, if it did then we could never generalize it. f(x) shifts left and right by adjust the value of x by some constant f(x-a) shifts the function to the right by a, why? let f(x) = k when x=0 f(x-a) = k when x=a x is shifted to the right by a

OpenStudy (amistre64):

x^2 + 1 = 1, when x=0 (x-4)^2 + 1 = 1 , when x=4 (x-4)^2 + 1 is the same function as x^2+1 shifted by +4

OpenStudy (amistre64):

cos(t) = 0 when t = pi/2 cos(t - pi/4) = 0 when t = pi/2 + pi/4 ^^^^^^ notice that the function is shifted to the right, by pi/4

OpenStudy (bloomlocke367):

okay